footer59:

Solve x2 + 6x + 7 = 0. x = −1 and x = −5 3 plus or minus square root of 2 negative 3 plus or minus square root of 2 quantity of negative 3 plus or minus square root of 2 all over 2

Salmon:

In the place of x put -1 and show me what you get

footer59:

i got -1 i think

Salmon:

Watch this

Salmon:

(-1)2 + 6(-1) + 7 = 0. -2 -6 + 7 = 0 -8 + 7 = 0 ? = 0 (-5)2 + 6(-5) + 7 = 0 -10 -30 + 7 = 0 -40 + 7 = 0 ?=0

Salmon:

Solve that and tell me if the question mark is equal to 0

footer59:

its nottt

Salmon:

Good What is the square root of 2?

footer59:

1. somethin im dumb im sorry

Salmon:

It's 4

footer59:

sorryyy math aint my strong suit

DBOY:

BRUHH ITS FINE

Salmon:

It's fine (3+ 4)2 + 6(3 + 4) + 7 = 0. (7)2 + 6(7) + 7 = 0 14 + 32 = 0 ? = 0 (3-4)2 + 6(3-4) + 7 = 0 -2 -6 + 7 = 0 -8 + 7 = 0 ?=0

Salmon:

Are they the same...the question marks....and do they equal 0?

footer59:

nopee

Salmon:

Good

Salmon:

So that leaves us with our obvious answer

Salmon:

OH WHOOPS I am really sorry about that we have one more! LOL

footer59:

awhh aha youre good! i was sittin there like well see sir its not too obvious to me quite yet lol

Salmon:

(-3+ 4)2 + 6(-3 + 4) + 7 = 0. (1)2 + 6(1) + 7 = 0 2 + 6 = 0 ? = 0 (-3-4)2 + 6(-3-4) + 7 = 0 (-7)2 -6(-7) + 7 = 0 -14 - 32 + 7 = 0 ?=0

footer59:

yeah that isnt it either

footer59:

so quantity of negative 3 plus or minus square root of 2 all over 2

Salmon:

Correct good Job!

footer59:

thank you so much man i really appreciate your help

dude:

Quick check: 1. The equation likely is \( x^2 + 6x + 7 = 0\), where the first x is squared. 2. The square root of 2 is *not* 4. It's irrational. 3. A quick solution to this is to use the quadratic equation (shown below) \(\dfrac{-b\pm\sqrt{b^2-4ac}}{2a}\) Where each value is defined by \(a^2+bx+c\) in the original equation